OpenStudy (anonymous):

Find the midpoint and distance between the points (-4,-10) and (-3,-1)

OpenStudy (aravindg):

midpoint is got by \[\dfrac{( x_1+x_2)}{2},\dfrac{(y_1+y_2)}{2}\]

OpenStudy (aravindg):

For distance use distance formula

OpenStudy (anonymous):

whats the distance formula?

OpenStudy (aravindg):

\[distance=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

OpenStudy (anonymous):

i got (-3.5, -10.5) and \[7\sqrt{10}units\]

OpenStudy (aravindg):

(-3.5, -5.5) is midpoint

OpenStudy (anonymous):

yes! your correct sorry

OpenStudy (aravindg):

-11/2=-5.5

OpenStudy (anonymous):

I meant the original equation to be (-3, and -11) not (-3, -1)

OpenStudy (aravindg):

I see

OpenStudy (aravindg):

Hope you can do it now :)

OpenStudy (aravindg):

click best response if my answer helped you :)